# Shifted Interpolation for Differential Privacy

### v1

@@ D.1 Details for Figure 1
These tables show that compared to our results (Theorem 4.5), the standard GDP Composition bound for NoisyCGD (Theorem 4.4) provides essentially no privacy. This is because that standard bound incurs a large privacy loss in each epoch (the step in which the adjacent datasets use different gradients), and this privacy leakage accumulates indefinitely—whereas our analysis captures the contractivity of the algorithm’s updates, which effectively ensures that previous gradient queries leak less privacy the longer ago they were performed. See §3 for a further discussion of this. Combined with the lossless conversion enabled by our $f$-DP analysis, our results also provide better privacy than the state-of-the-art RDP bounds.
@@ D.5 Numerical composition of subsampled GDP
Here we mention details on the numerical procedure for calculating an $(\varepsilon,\delta)$-DP bound from an $f$-DP bound of the form $f=C_{p}(G(\mu))^{\otimes t}$. This is used in §4.4. This formula appears in multiple settings of NoisySGD, with each $C_{p}(G(\mu))$ representing the $f$-DP of subsampled Gaussian mechanism. This conversion process is important for both notions: First, while the composition can be approximated by CLT (A.5), in practice the approximation is not enough to guarantee whether the algorithm achieves a given privacy budget, typically expressed in $(\varepsilon,\delta)$-DP. On the other hand, if one can obtain an accurate collection of different $(\varepsilon,\delta)$-DP bounds, it can be converted into an $f$-DP bound of comparable accuracy due to the duality between the two notions [DRS22, Proposition 5 & 6].

### v2

@@ D.1 Details for Figure 1
These tables show that compared to our results (Theorem 4.5), the standard GDP Composition bound for NoisyCGD (Theorem 4.4) provides essentially no privacy. This is because that standard bound incurs a large privacy loss in each epoch (at the step in which the adjacent datasets use different gradients), and this privacy leakage accumulates indefinitely—whereas our analysis captures the contractivity of the algorithm’s updates, which effectively ensures that previous gradient queries leak less privacy the longer ago they were performed. See §3 for a further discussion of this. Combined with the lossless conversion enabled by our $f$-DP analysis, our results also provide better privacy than the state-of-the-art RDP bounds.
@@ D.5 Numerical composition of subsampled GDP
Here we mention details on the numerical procedure for calculating an $(\varepsilon,\delta)$-DP bound from an $f$-DP bound of the form $f=C_{p}(G(\mu))^{\otimes t}$; this is used in §4.4. This formula appears in multiple settings of NoisySGD, with each $C_{p}(G(\mu))$ representing the $f$-DP of subsampled Gaussian mechanism. This conversion process is important for both notions: First, while the composition can be approximated by CLT (A.5), in practice the approximation is not enough to guarantee whether the algorithm achieves a given privacy budget, typically expressed in $(\varepsilon,\delta)$-DP. On the other hand, if one can obtain an accurate collection of different $(\varepsilon,\delta)$-DP bounds, it can be converted into an $f$-DP bound of comparable accuracy due to the duality between the two notions [DRS22, Proposition 5 & 6].